OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 22.

OpenStudy (anonymous):

replace x by 22 f(22)=-2*(22)+8

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

Put f(x) = 22 \[-2x + 8 = 22\] \[-2x = 14\] x = -7

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

wow i made a mistake its actually f(x)= 22 so it would be 22=-2x+8 so 22-8=-2x x=-7 sorry

OpenStudy (anonymous):

@Andreslvc that is value of f(x) given and not x..

OpenStudy (anonymous):

yes correction writted, didn't read the question completely sorry :D

OpenStudy (anonymous):

Be careful next time buddy..

OpenStudy (anonymous):

is it -7

OpenStudy (anonymous):

yes -7